OpenStudy (anonymous):

A phonograph record with diameter 12 in. rotates at 45 rev./min. Find the distance that a point on the rim travels in one minute.

OpenStudy (anonymous):

@CyberShadow @iPwnBunnies @mathmale @satellite73

OpenStudy (anonymous):

revolutions

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

turn turn turn

OpenStudy (anonymous):

Hey!!! Anybody home? @hello1213

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@amistre64

OpenStudy (confusionist):

I have no freakin idea. I just wanted to say you have a cool profile picture.

OpenStudy (anonymous):

oh haha thanks!

OpenStudy (amistre64):

what is the circumference?

OpenStudy (amistre64):

it travels 45 circumferences in 1 minute ....

OpenStudy (anonymous):

hmm isn't circumference the area around a circle?? :/

OpenStudy (amistre64):

the perimeter, but yeah

OpenStudy (amistre64):

\[\pi=\frac{C}{d}\] \[d~\pi=C\]

OpenStudy (anonymous):

okay so what do I plug in... I know d=12 right?

OpenStudy (amistre64):

d=12 yes, and pi is pi ... C = 12 pi inches, multiply by 45

OpenStudy (anonymous):

So 1696.5 ?

OpenStudy (amistre64):

45*12 pi = 540 pi which is about 1695.6 yeah

OpenStudy (anonymous):

it says to give answers to significant digits,

OpenStudy (amistre64):

i never did understand the significant digits stuff that well.

OpenStudy (anonymous):

so I think 1700... thanks for the help!

OpenStudy (amistre64):

good luck ;)